OpenStudy (anonymous):

Can someone explaine this problem to me please? Solve the inequality. Graph the solution set. 26 + 6b ≥ 2(3b + 4)

OpenStudy (anonymous):

Can you factor the left side. You can factor a 2 there.

OpenStudy (anonymous):

the left side of what

OpenStudy (anonymous):

Of the inequality. You can factor a 2 out of 26 + 6b to make it look like the right side.

OpenStudy (anonymous):

so would b be \[b \le1\frac{ 1 }{ 2 } or b \ge1\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

Step by step. What do you have when you factor a 2 out of 26 + 6b?

OpenStudy (anonymous):

\[2(3b + 13) \ge 2(3b + 4)\]

OpenStudy (anonymous):

wait whats that?

OpenStudy (anonymous):

We factored a 2 out of the left side. What happens when you FOIL out the left side? You get the original expression back. That is factoring. But now it looks like the right side.

OpenStudy (anonymous):

my book isn't much help at all, I'm sorry im so frustrating to help:\

OpenStudy (anonymous):

It's ok. Just follow me step by stop, don't go so far ahead. Do you see how that factored, though?

OpenStudy (anonymous):

There's another way to do this as well. FOIL out the right side of the inequality. IE: Distribute the 2.

OpenStudy (anonymous):

sorry i was busy for a sec

OpenStudy (anonymous):

So if you FOIL the right side, what do you get?

OpenStudy (anonymous):

how do you foil?

OpenStudy (anonymous):

Distribute.